Schalger corporation makes a product with the following standard costs.

Inputs------------------ Standard Quantity or hours Standard Price or rate

Direct Materials-------------- 7.8 Kilos $1.00 per kilo

Direct Labor------------------- 0.4 Hours $18.00 per hour

Variable Overhead----------- 0.4 Hours $3.00 per hour

The company reported the following results concerning this product in August.

Actual output----------------------------- 8500 units

Raw materials used in production--------------- 65,550 Kilos

Purchases of raw materials----------------- 69,000 kilos

Actual Direct labor-hours--------------- 3,410 hours

Actual cost of raw materials purchases---------- $75,900

Actual Direct Labor cost------------------ $ 66, 495

Actual variable overhead cost ----------- $ 9,889

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

Compute the materials quantity variance.

Compute the materials price variance.

Compute the labor efficiency variance.

Compute the direct labor rate variance.

Compute the variable overhead efficiency variance.

Compute the variable overhead rate variance.
Answer #1

a) CALACULATION OF MATERIAL QUANTITY VARIANCE:

Material quantity variance

=(Actual quantity−Standard quantity)×Standard Cost

=[65,550−(8,500×7.8)]×$1

=(65,550-66300)×$1

= 750 kilos (Favorable)

b) CALACULATION OF MATERIAL PROCE VARIANCE:

Material price variance

=(Standard rate−Actual rate)×Actual quantity

=($1- $1.10)×69,000

=$6,900 (Unfavorable)

Note: Actual rate is calculated by dividing actual cost of raw material purchased by actual quantity of raw material ($75,900/69,000) $1.10.

c) CALACULATION OF LABOUR EFFICIENCY VARIANCE:

Labor efficiency variance

=(Standard hour−Actual hour)×Standard rate

=(3,400 - 3,410)×$18

=$180 (Unfavorable)

Note: Standard hours = (8,500X0.4) = 3,400 hours.

d) CALACULATION OF LABOUR RATE VARIANCE:

Labor rate variance

=(Standard rate−Actual rate)×Actual labor hour

=($18 - $19.50)×3,410

=$5,115 (Unfavorable)

Note: Actual rate is $19.50 ($66,495/3,410 hour).

e) CALACULATION OF VARIABLE OVERHEAD EFFICIENCY VARIANCE:

Variable overhead efficiency variance

=(Standard hour−Actual hour)×Standard rate

=(3,400 - 3,410)× $3

=$30 (Unfavorable)

f) CALACULATION OF THE VARIABLE OVERHEAD RATE VARIANCE:

Variable overhead rate variance

=(Standard rate−Actual rate)×Actual hour

=($3 - $2.90) × 3,410

=$341(Favorable)

Note: Actual rate is $2.90 ($9,889/3,410 hours)

______×______

ALL THE BEST

HAPPY CHEGGING